Consider the following projects: Cash Flows (S) -1,400 2,800 3,500 1,400 1,400 1,400 1,400 1,000 1,400 1,400 4,400 1.400 1,400 a. If the opportunity cost of capital is 10%, which project(s) have a positive NPV? Positive NPV projectis) Project A Project B ProjectC Projects A and B Projects A and C O Projects B and C Projects A, B, and C No project b. Calculate the payback period for each project: (Round your answers to 2 decimal places. If a project never pays back, enter O.) Project A Project B Project C year(s) year(s) year(s) C. Which project(s) would a firm using the payback rule accept if the cutoff period were three years? Project(s) accepted d. Calculate the discounted payback for each project. (Do not round intermediate calculations. Round your answers to 2 decimal places. If a project never pays back, enter O.) (Click to select) * Projoct A Project B Project C year(s) years) year(s)
Answer #1

a) The Project B and C has positive NPV

Project A = -1400+1400*.909 = -127.40

Project B = -2800+1400*.909+1400*.826+4400*751+1400*.683+1400*.621 = 4759

Project C = -3500+1400*.909+1000*.826+0+1400*.683+1400*.621 = 424.40

b)

Project A 1 years
Project B 2 years
Project C 3.9 years

Payback period = Cash Inflow/ Cash Outflow

c) The firm would accept Project A and B if the cut off was 3 years

d)

Project A more than 5 years but exact cannot be calculated as cash flow is only in 1 Year and the discounted figure is less than 1400 years
Project B 2.1 years
Project C 4.6 years